OpenStudy (kymber):

Why doesn't 6÷2(1+2) equal 1?

OpenStudy (kymber):

Everywhere I look, people say it's 9. Shouldn't you follow PEMDAS? Parenthesis, 1+2 = 3 So then it's 6÷2(3) According to PEMDAS, you should do Multiplication and then Division, right? So shouldn't you do 2*3 = 6 And then 6÷6 = 1?

OpenStudy (anonymous):

You're right. lol

OpenStudy (anonymous):

Indeed you are right ,we can use bodmas also : B-Brackets first O-Orders (ie Powers and Square Roots, etc.) DM-Division and Multiplication (left-to-right) AS-Addition and Subtraction (left-to-right)

OpenStudy (anonymous):

PEMDAS is correct however, even in PEMDAS, you do multiplication and division from left to right. :P Same with addition and subtraction, left to right. So, the answer is indeed 1.

OpenStudy (anonymous):

I'm sorry, I mean it's 9. xD

OpenStudy (kymber):

Yeah, I forgot about the left-to-right thing! :p

OpenStudy (anonymous):

:)